OpenStudy (anonymous):

What is the correct classification of the system of equations below? 14x + 2y = 10 y + 7x = -5

OpenStudy (mathstudent55):

Divide the first equation by 2. Then write the x and y terms in the same order as in the second equation.

OpenStudy (mathstudent55):

What did you get?

OpenStudy (anonymous):

are they parallel?

OpenStudy (mathstudent55):

Did you do what I asked you to?

OpenStudy (anonymous):

yes i had that already i was asking what classification they were

OpenStudy (anonymous):

A. parallel B. coincident C. intersecting D. x = 10, y = 2

OpenStudy (mathstudent55):

If you did what I asked you to do, then now subtract one equation from the other. What do you get?

OpenStudy (anonymous):

wouldnt they just all cancel out?

OpenStudy (mathstudent55):

I don't know. I haven't solved this problem. I 've asked you to do a few things, and I haven't seen anything you've done, so I can't tell.

OpenStudy (anonymous):

y=-7x+5 y=-7x-5 they just cancel out if i subtract right?

OpenStudy (mathstudent55):

Ok. Now I see. Now subtract the second equation from the first equation. What is y - y?

OpenStudy (anonymous):

they cancel each other out....right? 0

OpenStudy (mathstudent55):

y = -7x + 5 y = -7x - 5 I just reread what you wrote above. You are correct. The lines are parallel. That means there is no solution to this system of equations.

OpenStudy (anonymous):

ty

OpenStudy (mathstudent55):

You can see the lines are parallel because they have the same slope and different y-intercepts.

OpenStudy (anonymous):

oh alright

OpenStudy (mathstudent55):

Also, if you subtract the equations, y = -7x + 5 (-) y = -7x - 5 -----------------(subtract) 0 = 0 + 10 0 = 10 you get 0 = 10 which is false. That means the system of equtions is inconsistent, meaning it has no solution.

OpenStudy (mathstudent55):

The reason I told you the above is that there are two ways of concluding this system has no solution. 1. The graphical way is: by noticing the lines a parallel. Since they don't intersect, there is no solution. 2. The algebra way: by noticing that when the equations are subtracted, you get a false statement, that means there is no solution.

OpenStudy (anonymous):

ok, thank you again for you in-depth explanations

OpenStudy (mathstudent55):

wlcm